OpenStudy (anonymous):

solve the quadratic equation. if no real numbers or roots state this.if involves radicals give both the radical form or the answer and calc. x^2-6x-2=0

OpenStudy (anonymous):

Using the discriminant, b^2-4ac, we can tell that this has 2 real roots because it is greater than 0.

OpenStudy (anonymous):

Use the quadratic equation to solve them

OpenStudy (anonymous):

\[\frac{6 \pm \sqrt{(-6)^2-4(1)(-2)}}{2(1)}\] \[\frac{6 \pm \sqrt{44}}{2}=\frac{6 \pm 2\sqrt{11}}{2}=3 \pm \sqrt{11}\]

OpenStudy (anonymous):

where did we get the 11

OpenStudy (anonymous):

\[\sqrt{44}=\sqrt{4*11}=2\sqrt{11}\]

OpenStudy (anonymous):

ok that makes sense